Negative divided by a positive. As of January 2007, the
national debt was $8.686 trillion dollars and the U.S.
population was 300.814 million people. Find the amount
of the debt per person to the nearest dollar.

Answer by Theo(13342)

8.686 trillion is equivalent to 8.686 * 10^12

thousand is 10^3
million is 10^6
billion is 10^9
trillion is 10^12

300.814 million is equivalent to 300.814 * 10^6

debt per person is 8.686 * 10^12 divided by 300.814 * 10^6

by the laws of exponents, 10^12 divided by 10^6 = 10^(12-6) = 10^6.

8.686 / 300.814 = .028874986

debt per person is therefore equal to .028874986 * 10^6.

This is equivalent to $28,874.99 per person.